OpenStudy (anonymous):

Help? What is the standard form equation of the line passing through (3, –2) with a slope of 3/5? This question is multiple choice and none of the answers are set up how I work this problem. Very confuzzling...:(

OpenStudy (anonymous):

Did you start from using the point slope form first because that's how i got my awnser

OpenStudy (anonymous):

Then the slope intercept the The standard form

OpenStudy (anonymous):

standard form is a^2+b^2=c^2?

OpenStudy (anonymous):

I keep coming up with something crazy like y=x-9.1....

OpenStudy (anonymous):

\[y+2=(3/5)(x-3)\]??

OpenStudy (anonymous):

\[y+2=(3x-9)(5x-15)\]??

OpenStudy (anonymous):

\[y+2=(3/5)(x-3)\] \[=(3x-9)(5x-15)\] \[=(15x-135)\] y=x-9.3...?